The Linux Distribution HOWTO

maintained by Eric S. Raymond v4.9, 10 December 1996 This document lists the various Linux distributions, repackagings, and services available for Linux via mail order, anonymous FTP, and other sources.

Contents

1 Introduction 2 1.1 New versions of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Recent Changes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Overview of the Linux Market . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Editorial Recommendations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 American Linux Distributions 5 2.1 Craftworks Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Debian Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 DILINUX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4 Linux Pro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5 Red Hat Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.6 Trans-Ameritech Linuxware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.7 Walnut Creek Slackware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.8 Yggdrasil Plug-and-Play Linux CD-ROM and the Linux Bible . . . . . . . . . . . . . . . . . . 13

3 International Distributions 15 3.1 S.u.S.E. GmbH German Linux CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2 Unifix 1.08 CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

4 Re-Packagers 18 4.1 InfoMagic Developer’s Resource CD-ROM kit . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.2 Oxbow Distribution Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

5 Miscellany 20 5.1 Unifix Custom Linux CD-ROM Mastering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.2 Linux Journal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.3 Sequoia International Motif Development Package . . . . . . . . . . . . . . . . . . . . . . . . 21 5.4 SSC Books and Linux Documentation Project Manuals . . . . . . . . . . . . . . . . . . . . . 22 5.5 Extent Verlag, LDP Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.6 UPython SLS and Slackware Diskette Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

6 Submissions To This Document 24 6.1 Types of Submissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2 How to submit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.3 Submission Guidelines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

7 Administrivia 26 7.1 Terms of Use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.2 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

1 Introduction

There is no single distribution of the Linux software. Instead, there are many such distributions, available both via anonymous FTP and mail order on floppy, tape, or CD-ROM. This document is an attempt to present information on many of the available distributions of Linux. This document also contains information on many miscellaneous services and goods available for Linux, ranging from consulting and support to T-shirts. The purpose of this document is to provide short summaries of the many Linux distributions and mail-order services, and to provide pointers for the reader to find more information. The information presented here is far from complete; there are many more Linux distributions and services than are listed here. Unfortunately, we have not received submissions from many of the organizations providing these services. Please see 6 near the end of this document for information on making a submission. It’s easy and will take less then five minutes. Disclaimer: We make absolutely no guarantee as to the correctness of the information, prices, and ordering details given in this document. Check the last-modified field of each to get an idea of its currency. Furthermore, unless otherwise stated the Linux software comes with ABSOLUTELY NO WARRANTY . Your editor tries to stick to facts in most of this HOWTO, but he has some opinions on the state of the Linux market. If you care what they are, you can read them under 1.4. Disclosure: I (esr) have no financial connection to any Linux vendor, nor have I accepted any renumeration or perquisites from any vendor other than free product for review (and one T-shirt from Red Hat).

1.1 New versions of this document

This document will be posted monthly to the newsgroups comp.os.linux.answers . The document is archived on a number of Linux FTP sites, including sunsite.unc.edu in pub/Linux/docs/HOWTO.

You can also view the latest version of this on the World Wide Web via the URL

<http://sunsite.unc.edu/mdw/HOWTO/Distribution-HOWTO.html>.

Feel free to mail any questions or comments about this HOWTO to Eric S. Raymond, esr@snark.thyrsus.com. Please do not send me general Linux questions or requests for help in choosing a distribution unless you’re willing to hire me at normal consulting rates; I don’t have time to deal with them, and I try to put everything I know about choosing a distribution in this document.

1.2 Recent Changes

Network distributions like the original Slackware no longer seem to be Linux’s most important vector. Accordingly the General Information section and this HOWTO as a whole now focuses more on commercial CD distributions.

I have dropped the section on Linux mail-order resellers (Spheric Microsystems, Takelap, Lasermoon). You can’t turn around in a computer store, computer hobbyist fair, or even large mainstream bookstore any more without tripping over a pile of Linux books and CD-ROMs available cheap. So tracking these resellers is no longer a service to readers worth the work it costs me.

NOTE: I am planning on dropping the section on 3 for reasons discussed there. If you are appropriately multilingual and are concerned about the quality of that section’s information, please volunteer either to (a) maintain a International Distributions HOWTO, or (b) act as part of my international spy network, sending me regular updates I can use.

1.3 Overview of the Linux Market

In the beginning (say, 1993), a Linux distribution was something you downloaded off the Internet onto floppies. Installation was a laborious process and repeated frustrations due to bad media were common.

Then came cheap CD-ROM drives and the CD-ROM, a medium ideally suited for shipping large volumes of operating-system software cheaply. There’s a whole mini-industry now built around commercial CD-ROM Linuxes, and (because the vendors have actual cash flow to fund support and marketing) they increasingly dominate the Linux world. Debian is now the only significant non-commercial release, and even it seems to be propagated largely by shovelware CD-ROMs.

Most of the CD-ROM distributions (including Slackware, Yggdrasil and Red Hat) are still available for FTP from the home sites of their developers. But if you have a CD-ROM drive and a few dollars, you will have many more distributions and more support options to choose from (and you’ll usually get some useful paper documentation). For more on the details of installation, see the Linux Installation HOWTO,

<http://sunsite.unc.edu/mdw/HOWTO/Installation-HOWTO.html>.

There are three tiers in the Linux CD-ROM market. The top tier consists of primary distribution builders (such as Yggdrasil, Red Hat, WGS, and Craftworks). These people add substantial value to Linux, often including custom administration/installation tools, bound documentation, and availability of technical support contracts.

The bottom tier consists or redistributors and publishers who simply shovel Slackware or some other network-available distribution onto a CD-ROM and re-sell it with zero or minimal documentation and support.

The middle tier is harder to define; it consists of re-packagers who add some value without actually building and maintaining their own primary distributions.

One thing that does not distinguish these tiers much is price. Prices for CD-ROM distributions of Intel Linuxes start at $20 and top out at a whole $50 (and the extra few dollars can buy real value). Many vendors sell subscription deals that will lower your cost-per-CD for regular updates over the subscription period.

In October 1996 price correlates with features and quality pretty well (as one would expect in a very competitive market). Your editor recommends paying the few extra dollars for a top-drawer CD-ROM distribution; this will pay off in fewer installation and administration hassles down the road.

1.4 Editorial Recommendations

Last section, the facts. In this section, my opinions (for whatever they’re worth – and remember the caveat about free advice). There is no substitute for doing your own evaluation based on experience and the data in this guide, and these are intended more to illuminate my possible biases than as a guide to what you should do.

From the beginnings of the Linux CD-ROM industry in 1993 to Fall 1995, Yggdrasil was the king of the hill – it essentially founded the CD-ROM market and then set the standard for everybody else. I used Yggdrasil, and I recommended over commercial System V versions for their superior documentation, large collection of applications, and enlightened policy of sending free releases to freeware authors and dedicating part of the price of each CD-ROM to financially supporting free software. But Yggdrasil hasn’t issued a new release in 1996 (it’s December as I write) and they’ve perhaps been left behind by the market.

I now run Red Hat Linux and am quite satisfied with it. Red Hat’s RPM technology currently gives it, IMO, a technical edge over any other vendor. They’ve made most of the right moves at the right times and I consider them the current market leader.

However, these opinions should not be interpreted as an unconditional endorsement; different Linux distributions are optimized for different needs, and yours may well be best served by some other distribution (especially if, unlike me, you’re mainly a DOS user and are looking for a distribution tuned for dual-boot systems and being launched from DOS).

Furthermore, industry standing is volatile. By the time you read this, Red Hat may well have fallen off its game and been displaced by a hungrier newcomer.

2 American Linux Distributions

All these distributions are available on CD-ROM (some of the non-commercial ones only show up on network-archive snapshots). Most of these are available for free over the network (but the commercial ones won’t support you if you buy this way). They are custom-assembled, rather than just being re-packagings of a pre-existing network release.

They are listed in alphabetical order.

2.1 Craftworks Linux

Distributor:

Craftwork Solutions, Inc. 4320 Stevens Creek Blvd. #170 San Jose, CA 95129 Net: info@craftwork.com WWW: http:/www.craftwork.com Vox: (408)-985-1878 Fax: (408)-985-1880

Provider’s Description:

Craftworks Linux v2.2 is a commercial distribution of Linux for Intel and AXP. It is completely ELF, built one program at the time from the latest ”stable” sources and organized using a methodology we call: CRAFT (Component Replacement And Fabrication Technology). CRAFT accounts for each file in the system, groups them logically as true components which are filed into catalogs. For example from the ’System Install’ catalog, the ’Main Operating System Component’ which accounts for 3788 files (approx. 19MB), when combined with the ’Linux Standard Devices’ component constitute an operational, network ready, GNU/Linux system. CRAFT can now install RPM packages. Security measures including shadow password are enforced throughout the distribution. Network information services (NYS) and domain name services (DNS) as provided, coexist to complement each other. Screen savers and a desktop environment are there for X Windows. Extensive use of kernel loadable modules including streaming tape (ftape) support have been utilized to minimize the need to provide various kernel configurations or requiring the end user to build one. These are some of the highlights, and the overall thought that went into this distribution, which provides a powerful modern UN*X environment for workstations and servers. The Craftworks Linux v2.2 package includes:

• A boot/install floppy for installing on one or across multiple hard-disks from the CD-ROM media or via NFS (ethernet or PPP). Craftworks Genesis guides the user with panels and menus through a new install or upgrade.

• One CD-ROM with a live system (i.e. run directly from the CD-ROM) and all the sources as GNU zipped tar files. Supporting files such as a comprehensive list of CREDITS, the GNU Public License, etc. are also found on the CD-ROM media and copied as part of an install onto the hard-disk. Craftworks Sysadm (Main Operating System Component) provides for:

– Security and User Account Management

– System Maintenance (access to catalogs)

– BRU Personanal Edition backup utilities

– X Window System Setup including Metro X enhanced server.

– UPS Configuration

– System Setup

• Reference and Installation Guide with chapters covering the installation and maintenance thereafter. Appendices on FAQs and hardware compatibility.

Internet Access:

Updates between releases are found in <ftp://ftp.craftwork.com/pub/v2.2>. FAQs and up-to-date product information <http://www.craftwork.com>.

Ordering:

Directly from Craftwork Solutions, Inc. or from distributors and stores. Suggested retail price 59.95(Intel),89.95 (AXP). Intel (80386/486/Pentium) version. Discounts available to dealers and educational institutions.

Support:

Craftwork Solutions, Inc. provides 30 days support with each copy. Longer term support as well as special projects support available on a contract or fee basis.

Miscellaneous:

The CRAFT-1.0 specification is available from <http://www.craftwork.com>. It includes a complete component listing for the distribution.

Last Freeze Date:

23 August 1996.

Entry last modified:

23 September 1996.

2.2 Debian Linux Distribution

Distributor:

The Debian Linux Association Software in the Public Interest P.O. Box 70152 Pt. Richmond CA 94807-0152 Net: info@debian.org

WWW: <http://www.debian.org/>
FTP: <ftp://ftp.debian.org/debian/>

Provider’s Description:

Debian Linux is the result of a volunteer effort to create a high-quality non-commercial Linux distribution. Debian Linux 1.1 is a complete and fully-functional Unix-compatible operating system for the personal computer. The system uses ELF executables, but also includes optional packages you can install to fully support development and maintenance of a.out programs. At present, Debian Linux is available for the IBM PC architecture (386/486/Pentium) with versions for Sparc, m68k and Alpha architectures well under way. Debian Linux is an easy-to-upgrade distribution that mostly consists of freely redistributable software from a number of different sources and systems. Support for shadow passwords is available, although it should be considered experimental in the 1.1 release. Non-free packages may also obtained, making Debian a well-rounded system. The benefits of Debian Linux are its upgradability, well-defined dependencies between packages, and its open development. It is the only distribution of Linux that is being developed co-operatively by many individuals through the Internet, in the same spirit as Linux and other Free Software. More than 100 package maintainers are working on about 500 packages and improving Debian Linux. A sophisticated bug tracking system allows users to easily report bugs and security concerns which are quickly dealt with by the Debian community. A new release of the package will soon occur on well-known ftp archives. Debian Linux is a very dynamic distribution. Snap-shot releases are made about every three months, the ftp archives are updated daily. For more information about Debian Linux, please refer to the files at

<ftp://ftp.debian.org/debian/doc/> or visit our World Wide Web page at
<http://www.debian.org/>.

If you’re interested in joining this project you are invited to subscribe to either debian-user@lists.debian.org or debian-devel@lists.debian.org by sending a mail to debian-user-request@lists.debian.org (or debian-devel-request respectively) with the word ”help” in the subject line.

Internet Access:

Debian Linux is available via anonymous FTP from <ftp://ftp.debian.org/debian/> and on a mirror near to you.

Last Freeze Date:

27 July 1996

Entry last modified:

8 October 1996

2.3 DILINUX

Distributor:

Kent Robotti

FTP: <ftp://ftp.sunsite.unc.edu/pub/Linux/distributions/dilinux>

Provider’s Description:

A (D)rop (I)n linux-elf slip/ppp networking system. It can be dropped into a subdirectory of any DOS system and booted from DOS without messing with disk partitions. About 22 mbytes unzipped, 8 1/2 mbytes zipped.

Internet Access:

By anonymous ftp from Sunsite archives (see URL above).

Last Freeze Date:

30 Sep 1996

Entry last modified:

8 Oct 1996

Editor’s comments:

As the author says, not a general-purpose release. Probably best suited to sites that primarily run DOS but want better networking tools for hooking up to an Internet service provider.

2.4 Linux Pro

Distributor:

WorkGroup Solutions, Inc. P.O. Box 460190 Aurora, CO 80046-0190 Vox: (303)-699-7470 Fax: (303)-699-2793 Net: info@wgs.com (orders)

FTP: <ftp://ftp.wgs.com/pub2/wgs>

Provider’s Description:

The charter of the WGS Linux Pro CD is different from all other distributions currently on the market. We are actively pursuing penetration into the commercial marketplace for Linux. Everything we do is aimed at increasing the size of the Linux community, and helping Linux to become THE one true operating system, not just an operating system. WGS Linux Pro consists of our Main ”Stable” Linux CD, for which we select what we consider to be the best Linux distribution available. Then we make bug fixes & minor enhancements. This CD often contains software older than the latest, under the theory that ”the latest is not always the greatest”. Together with this approach we provide technical support as well as support agreements. For this reason WGS Linux Pro is considered a primary Linux distribution. Optionally, and in addition to our primary CD, we supply supplemental CDs containing all the latest Linux software (including current archives of the tsx-11, sunsite, and Red Hat sites). Virtually anything you will get with anyone’s Linux product can be found on these supplemental CDs. We continually add more to these supplemental CDs as software becomes available. Even with all this, pricing on our product line is very favorable when correctly compared to other Linux distributions. Also optional (included with Linux Pro+) is a professionally produced and printed manual containing the Linux documentation project and more! Current version is 4.0a; kernel is 1.2.13 with installable 2.0.x on the CD.

Internet Access:

Look in <http://www.wgs.com> for a list of WGS products, and latest information on the CD-ROM.

Ordering:

WGS Linux Pro is available on CD from dealers and distributors worldwide, and soon from our FTP site. You may call, fax, email, or mail us, to purchase, or request additional information. We will be happy to respond to any question you have. We accept American Express, Visa, Discover, and MasterCard, or purchase orders with approval of credit.

Miscellaneous:

Complete information is available on our FTP site. WGS publishes a free E-Mail newsletter on Linux and FlagShip to which you may subscribe by just emailing us your request. In addition we carry Motif, FlagShip, A line of books, and other products to make your experience with Linux more complete and fulfilling. Check us out!

Last Freeze Date:

1 Oct 1996.

Entry last modified:

8 Oct 1995.

Editor’s comment:

They describe their current distribution as ”Red Hat 3.03 plus”.

2.5 Red Hat Linux

Distributor:

Red Hat Software 3201 Yorktown Rd, Suite 123 DeKalb Center Durham, NC 27713 Vox: (800) 546-7274 or (919) 572-6500 Fax: (919) 572-6726 Net: redhat@redhat.com

WWW: <http://www.redhat.com>
FTP: <ftp://ftp.redhat.com>

Provider’s Description:

Red Hat 4.0 (Colgate) is based on the 2.0.18 kernel and is available for Intel, Sparc and Alpha platforms. Four installation innovations make Red Hat the easiest Linux to install ever. Our graphical installation mode brings you straight up into X for most of the installation. Our boot disk creation script eases selection of the proper boot disk, and it saves your existing network configuration information and XF86Config so you don’t have to configure TCP/IP or X! Our FTP install allows you to install simply by downloading 3 floppy disk images – the rest is done automatically! And finally, the installation includes seamless support for PCMCIA devices – install Red Hat on your laptop as easily as on your desktop machine! After installing Red Hat Linux once, you will never need to reinstall Linux again! The new RPM packaging system is sophisticated enough to allow upgrading to new Red Hat releases without reinstalling your system - no partitioning, no backing up all your files, no headaches.

Ordering:

Available directly from Red Hat Software and most distributors of Linux related products. Please contact Red Hat Software for ordering details (on-line ordering is available through Red Hat’s Web). Suggested retail price is for the Intel version is $49.95, for the Sparc and Alpha versions $99.95.

Last Freeze Date:

5 September 1996.

Entry last modified:

23 September 1996.

Editor’s Comments:

The big selling point of this distribution is RPM, the Red Hat Package Manager. This piece of software is a remarkable advance; it allows you to cleanly install and de-install applications and operating-system components, including the kernel and OS base itself. Red Hat has issued this under GPL, which is a good thing for other distributors; it is the considered opinion of this editor that those who do not adopt it will shortly have their lunches eaten by those who do.

2.6 Trans-Ameritech Linuxware

Distributor:

Trans-Ameritech 2342A Walsh Avenue Santa Clara, CA 95051 Net: info@trans-am.com order@trans-am.com WWW: http://www.zoom.com/tae Vox: (408)-727-3883 Fax: (408)-727-3882 BBS: (408)980-9840 Provider’s Description:

Trans-Ameritech has published 8 releases of Linux by the end of 95. Traditionally, Trans-Ameritech has set new standards for combining ease of use and straightforward installation, even for a first-time user, with its line of LinuxWare CD-ROMs. This is a flexible, easy-to-install operating system geared toward those intersted in learning Unix as well as technical people, students and home PC users.

• New Linux users will appreciate a Windows-based set up program from the CD.

• To minimize the possibility of hardware conflicts many extra kernels are provided for different configurations. They are usable for installation and normal operation.

• Many on-line documents are provided for quick reference, including the Linux Documentation Project files in source, dvi and ps formats.

Many applications are included:

• C/C++,Pascal and Ada compilers as well as converters fron Fortran

• TCP/IP networking, UUCP,SLIP,CSLIP

• Internet access with binaries and sources for FTP,Telnet,News and E-mail

• Multiple terminals and X-Windows environment

• Gnu and international versions of the ispell spell-cheker

• The communications apps:term,minicom,Seyon ( X-Windows based)

• Editors:elvis(vi clone),joe,jove

• PostScript clone ghostscript

• Object oriented GNU Smalltalk and the Smalltalk interface to X

• TCL/Tk (Powerful scripting language with Motif-like X interface)

• Programs for electrical engineers and HAM enthusiasts

• Windows based X configuration program ???

• Interviews libraries,include files and doc Word-processor and idraw drawing program

• Typesetting:TeX,LaTeX,xdvi,dvips,Metafont,groff

• Andrew multimedia word prodessor with hyperlinks

• FAX send and receive on either class 1 or class 2 fax modems

• DOOM for game enthusiasts

All the sources are available on the CD-ROM. The often needed sources are uncompressed and can be used directly from the CD-ROM. An uncompressed Linux filesystem is available for references and disk space conversation. You can run programs directly from the CD-ROM! There is a large info directory for on-line reference and many manpages. For hacker’s reference, uncompressed 386 BSD sources are provided on Supplement 4. The latest upgrades and patches are always available through Trans-Ameritech Linux BBS:(408) 980 9840. It’s free for all Trans-Ameritech customers. Our distribution is targeted for Windows and DOS users who want an easy migration path upwards,rather than for UNIX experts.

Ordering:

The price for our current release of LinuxWare 2.5 is $19.95 The price for a package order: LinuxWare 2.5,Supplement 3 and Supplement 4 (Ultimate LinuxWare Bundle) in one shipment is $30 The book ”Running Linux”- a 550 page comprehensive Linux tutorial by Matt Welsh is $20 If you order with a credit card (VISA, MC, AmEx, Discovery), please, indicate the card number,expiration date and your mailing address. Shipping and handling in US: single CD-ROM or Ultimate LinuxWare Bundle is $5(first class US mail).If you would like to order CD-ROM with the book,shipping is $8(UPS ground). Overseas shipping is $8 for single CD-ROM , $12 for the bundle and $25 for the bundle plus the book. COD is available in the US only for $4.50. California residents, please add 7.75% sales tax. Annual subscription (4 releases) are available for $80 plus S&H (note: there are 4 shipments in a subscription). Example: subscription in US is:$80+$5x4=$100. Subscription in Europe/Japan etc. $80 + $8x4 = $112.

Miscellaneous:

If you have any further questions,please contact us through E-mail: info@trans-am.com or order@trans-am.com.

Last Freeze Date:

September 1995 Entry last modified:

29 November 1995.

2.7 Walnut Creek Slackware

Distributor:

Walnut Creek CDROM 4041 Pike Lane, Suite D Concord, CA 94520 Net: info@cdrom.com (information), order@cdrom.com (orders), support@cdrom.com (support).

Provider’s Description:

Ftp.cdrom.com is the home of Slackware Linux. We are the publishers of the Official Slackware Linux CDROM. Our distribution is a 2-disc set with the current version being Slackware 3.0. The current disc contains kernels 1.2.13 and 1.3.18 as well as Elf Binaries. Slackware Linux is a full featured distribution of the Linux operating system designed for 386/486 computers with a 3.5" floppy and CD-ROM drive.

Internet Access:

WWW: <http://www.cdrom.com/titles/slackware.html>
FTP: <ftp.cdrom.com/pub/slackware>

Ordering:

The cost of the CD-ROM set from Walnut Creek is $39.95.

Editor’s Comments:

A lot of second- and third-tier commercial Linuxes are re-packagings of Slackware (see the Re-Packagers section below.

Last Freeze Date:

September 1995.

Entry last modified:

23 November 1995.

2.8 Yggdrasil Plug-and-Play Linux CD-ROM and the Linux Bible

Distributor:

Yggdrasil Computing, Incorporated 4880 Stevens Creek Blvd., Suite 205 San Jose, CA 95129-1034 Net: info@yggdrasil.com

WWW: <http://www.yggdrasil.com>
FTP: <http://ftp.yggdrasil.com>

Vox: (800) 261-6630, (408) 261-6630 Fax: (408) 261-6631

Provider’s Description:

Yggdrasil Plug-and-Play Linux is a complete CD-ROM distribution of the Linux operating system. It includes a great deal of software—nearly every package that you would expect to find on a complete UNIX system is available. A complete file list is available via FTP from yggdrasil.com. The Linux Bible is a compendium of Linux documentation, including 5 books from the Linux Documentation Project, The Yggdrasil installation manual, and the complete set of Linux HOWTO guides. We publish this and offer this separately from the CD-ROM distribution. At the top of the sophistication hierarchy, programmers who want to explore or add the occasional feature know that Yggdrasil is the Linux distribution with a fully buildable source tree and with the ability to automatically trace installed files back to their sources. Everybody, especially new users, will appreciate the Plug-and-Play operation for which Yggdrasil’s product is named. Put the media in a computer with supported hardware, turn the computer on, and it’s running everything, straight from the CDROM. (This is no longer unique to Yggdrasil, but Yggdrasil did it first.) The login screen lists a number of preconfigured user names, including ”install”, which installs the system, giving paragraphs of explanation about every question that it asks the user. X windows configuration is automated too, prompting the user for configuration information the first time xinit is run.

From X windows, a graphical control panel allows simple “fill in the blanks” configuration of networking, SLIP, outgoing UUCP, the printer, NNTP, and many other features that previously required the knowledge of a system administrator to configure. For more information, send mail to info@yggdrasil.com, surf to www.yggdrasil.com, FTP to yggdrasil.com, or contact us by any convenient method.

Ordering:

Plug-and-Play Linux costs $34.95 is and available directly from Yggdrasil or from your local computer, software or technical book store. If Plug-and-Play Linux is not available from your favorite reseller, help promote Linux by making it your mission to change that. Give your reseller our phone number and demand that they carry Plug-and-Play Linux. Yggdrasil offers a $7 discount for upgrades or crossgrades. Send us your old Yggdrasil release and a check for $27.95 + $5 shipping and handling to upgrade. Or do the same, but send us a competing distribution such any version of SCO, Esix, minix, or one of the CD’s with the slackware floppy images, and tell us where you got it, so that we can make sure that your favorite reseller carries our products too.

Miscellaneous:

Yggdrasil also sells OSF/Motif and The Linux Bible, a compilation of works from the Linux Documentation project. In addition to a copy of the Plug-and-Play Linux manual, The Linux Bible includes Linux Installation and Getting Started, Network Administrator’s Guide, System Administrator’s Guide, Linux User’s Guide, and HOWTO guides on hardware compatibility, distributions, dos emulation, ethernet, floppy tape, installation, mail, networking version 2, MGR (an alternative to X windows), X windows, usenet news, printing, SCSI, serial communications, sound, and UUCP. The Linux Bible costs $39.95 and is printed on recycled paper. $1 from every copy sold is donated to the Linux Documentation Project. OSF/Motif costs $149.95, $5 of which is donated to the development of a free Motif clone.

Last Freeze Date:

Mid-August 1995. (Yes, 1995. They have a next release in pre-production but don’t want to commit to a ship date yet.)

Entry last modified:

10 December 1996

Editor’s Comments:

Yggdrasil includes lots of well-chosen stuff with their distribution; the CD-ROM-only demo mode is impressive; installation and administration are easy enough to live up to their ”Plug And Play” name. The 167-page manual is very nice. Their previous problems in the installation scripts seem to have been fixed. Their only drawback is that they haven’t yet adopted Red Hat’s RPM technology or created an equivalent. They claim they’re going to by RPM-compatible in their upcoming release.

3 International Distributions

These are distributions intended for use by non-English-speakers. So far, all happen to be German.

NOTICE: I am planning to drop this section soon! These suppliers have sent me documentation I’m linguistically incompetent to read. I can’t get fresh information on them in correct form for this HOWTO without making transatlantic phone calls. If you are appropriately multilingual and are concerned about the quality of this section’s information, please volunteer either to (a) maintain a International Distributions HOWTO, or (b) act as my part of my international spy network, sending me regular updates I can use.

3.1 S.u.S.E. GmbH German Linux CD-ROM

Distributor:

S.u.S.E. GmbH Gebhardtstrasse 2 90762 Fuerth Germany Net: bestellung@suse.de Vox: +49 911-74053-31 Fax: +49 911-7417755

Provider’s Description:

This is a (triple) CD mainly for German users. S.u.S.E. now presents their own distribution (starting in may 1996) indepent from slackware or other distributions. This is meant to guarantee a perfect interaction between all the packages on the CD and facilitates the menu-driven Installation and high conformity to the Linux filesystem standard. S.u.S.E. Linux contains additional software, such as the Postgres system, the pbm-Tools and lots of other tools and utilities (some window managers, GREAT environment, the POV- ray raytracer, a fully working and licensed database system (LNX) and so on). S.u.S.E. Linux also contains a couple of demo versions, preconfigured, such as OSS (Open Sound System), GraffyDOC (CAD), Maple,WordPerfect, YardSQL, Xess3. The source code of the complete system is included, too, as is a live system to enable using software without the need to install everything to the harddrive and not even partitioning. The current version (November ’96) of the distribution is 4.4. incorporating the 2.0.25 kernel. A printed German installation guide will help especially new users to successfully install the system. Since version 4.3 there is also an English version including English installation and an English manual. A 60 day customer support is included to help in serious installation crises ;-) via email, fax, snail mail or phone hotline (see below). There are also interesting bundles of S.u.S.E. Linux together with Metrolink Motif Runtime for DM 174.- (or DM 159.- Update to former S.u.S.E. Linux versions). Another product is S.u.S.E. Linux Aktuell!, a four CD ftp-server copy which appears about every 2 months. It contains the linux directory trees of the most popular ftp-servers such as SUNSITE.UNC.EDU, TSX- 11.MIT.EDU, FTP.FUNET.FI, PREP.AI.MIT.EDU (GNU Archiv), ftp.neosoft.com (TCL), ftp.phil.uni-sb.de (68k Linux), ftp.uni- erlangen.de (68k Linux), ftp.x.org (contrib-dir), elisp-Archive and RFCs. A basic S.u.S.E. Linux system is included, but no printed manual or customers support. New is ”Linux Decathlon”, an edition similar to Linux Aktuell! but containing 10! CDs with almost any ftp-server important for Linux including also general non-Linux ftp-servers.

Ordering:

By mail, see contact information above. Price: The price for a single German Linux CD is 98,– DM for new customers. If you are already a S.u.S.E.-customer, you will get the CD for an update-price of 79,– DM. There is also the posibility to subscribe for 4 subsequent editions; then you will receive 4 CDs for the price of 259,– DM. The ”Linux Aktuell!” CD is 45.- DM. The ”Linux Decathlon” CD is DM 89.–. Disks: Of course, you can get the German version of the Slackware Distribution on 3,5" disks. You can get either a basis version of the distribution (about 28 disks) for 89,– DM, the ”standard” version (about 38 disks) for 119,– DM or the ”full” version (about 59 disks) for 159,– DM. The update prices are 69,– DM, 94,– DM and 139,– DM respectively. Documentation: S.u.S.E. is selling high-quality printed issues of the books from the Linux Documentation project (LDP) for people that would like to have a real book and not only printed loose papers. The prices are 59,– DM for Matt Welsh’s Linux Installation and Getting Started, and 50,– DM for Olaf Kirch’s Network Administrators’ Guide. Additionally, S.u.S.E. offers the complete O’Reilly series which is the best documentation available for Unix in general. Of course you can order the well known German Linux books: Linux - Wegweiser fur Netzwerker; Olaf Kirch; DM 59,- Linux - Wegweiser zur Installation & Konfiguration; Matt Welsh & Lar Kaufmann; DM 59,- Motif: Metrolink Motif for Linux is available for the price of 99,– DM for the runtime system only and for the price of 239,– DM you will get the runtime and development system.

Support:

Service and support for customers by Email, Fax, snail mail (every day) and phone (hotline Monday and Thursday from 13:00h-18:00h.) (email: support@suse.de) 24h support database on WWW helps solving common problems with Linux and S.u.S.E. Linux: <http:/www.suse.de/Support/sdb> in

German or <http://Riemann.suse.de/Support/sdb e/index.html> in English.

Miscellaneous:

Free info material and catalogues of CDROMs and books. (All Infomagic Unix CDs available). Selling commercial software for Linux. Development of individual software for Linux and other Unix systems. Professional support for Linux and any other software (psupport@suse.de <mail:psupport@suse.de>). Almost anything concerning Linux - just ask!

Entry last modified:

18 November 1996

3.2 Unifix 1.08 CD-ROM

Distributor:

Unifix Software GmbH Bueltenweg 27a D-38106 Braunschweig, Germany Vox: +49 (0)531 238800 Fax: +49 (0)531 338751 Net: support@unifix.de

Provider’s Description:

POSIX.1 certified version 1.8 will be available at mid february 96. This is a Linux CD-ROM distribution with emphasis on easy and fast installation. Though it is possible to install everything on the hard disk, we do not recommend doing so, because it runs fast enough directly from CD. For example starting emacs the first time (in text mode) takes:

                From single speed Mitsumi: 24 s
                From double speed Mitsumi: 11 s
                From Toshiba 3401:          7 s

The dynamic buffer cache of Linux takes care of the second and subsequent times a program is started, which needs less than one second. Our system requires about 5 MB on the hard disk for configuration files and system management tools. Additionally we recommend a 16 MB swap partition. The distribution contains most of the standard programs in current versions, e.g. Linux 1.2.3, XFree 3.1.1, emacs 19.30 and gcc 2.7.2. Full preconfigured source for everything is included. Because Unifix is a European distribution it provides full support for iso8859-1 character sets. From the shells and emacs through ls, TeX and the print system everything supports 8-bit characters. Printing is supported through System V compatible printer drivers which can be controlled through lp’s -o options. Printer drivers are included for dumb text mode printers, for postscript- or ghostscript- compatible printers and for networked printing. These drivers know how to guess the type of the files so e.g. compressed man pages or DVI files can be printed directly.

Ordering:

Two versions are available: a small one called Quickstart and a POSIX.1 certified version with OSF/Motif (called developers release). Unifix is available only on CD. It comes in a Unifix/Linux binder with two boot floppies and about 70 pages of installation instructions in German. An English version is available from Lasermoon (Email support@lasermoon.co.uk). The price is 99 DM for Quickstart and DM 298 for the developer’s version, and includes taxes and shipping; Eurocard/Mastercard/Visa are accepted.

Entry last modified:

4 May 1996

4 Re-Packagers

This section lists re-packagers of the Linux distributions listed above. In other words, the people selling the software below probably do not maintain or support the software itself.

4.1 InfoMagic Developer’s Resource CD-ROM kit

Distributor:

InfoMagic, Inc. PO Box 30370. Flagstaff, AZ 86003-0370 Net: info@infomagic.com

WWW: <http://www.infomagic.com>

Vox: (800)-800-6613 (orders), (520)-526-9852 (tech support) Fax: (520)-526-9573

Provider’s Description:

The InfoMagic Linux Developer’s Resource is a snapshot of the sunsite.unc.edu and tsx-11.mit.edu archives. It also includes the complete GNU software collection (in source form). The following Linux distributions are included on the discs: RedHat, Slackware, Debian and JE (Japanese Extensions). The Slackware distribution has been completely unpacked allowing many packages to be run directly from the disk. Sources for all the packages in Slackware are also included. The Linux HOWTO documents have been formatted for use with the Microsoft Multimedia Viewer (which is included) to allow browsing and full-text search under Microsoft Windows. InfoMagic also includes complete archives of the linux mailing lists. InfoMagic also distributes the ”Installation and Getting Started Guide”, by Matt Welsh as well as Linux T-Shirts (featuring the platypus logo).

Ordering:

The 5-CD set is $25/copy. Shipping within the US is $5 (USPS Priority Mail), outside the US $10 (International Airmail), FedEx and UPS on request. Orders may be placed via phone, fax, or email (a PGP key is available: finger info@InfoMagic.com). We accept Visa, Mastercard, and AMEX. 1-year subscriptions are available for $150 (within the US) and $175 (outside the US) including shipping. A 1 year subscription consists of 6 releases, one every two months or so.

Internet Access:

The contents of the CDs may be found at either <http://www.infomagic.com>,

<ftp://ftp.infomagic.com>.

Support:

We provide unlimited email and FAX support. Telephone support is available via a 900 number at $2.00/minute. We also offer support contracts tailored to individual needs. The CD-set includes a 30-page quick-start.

Last Freeze Date:

30 November 1995. Updates about every two months. The Slackware release is 3.0, RedHat is 2.1 Entry last modified:

25 November 1995.

4.2 Oxbow Distribution Services

Distributor:

Oxbow Enterprises Net: pdykes@sparc.isl.net (orders) Vox: 507-775-7096 (1-800 service coming soon!)

Provider’s Description:

Our ”Beginner Pack”: Red Hat Linux CD(s), Dr. Linux (1200 pages of great info!) - $79.95 The Beginner Pack is $39.95 w/o Dr. Linux pages book (update customers get a discount, so cyou ould get it later at reduced rate). Linux Man Pages book to accompany the Beginner’s Pack is $41.95 (by itself, $49.95). We will do machine integration. We’ll integrate an Acer, Packard Bell, IBM, etc... with Linux and software purchasable with platform for cost of machine, shipping, + $139.95. Our objective is to offer Linux and Windows ’95 on brand name systems, providing new and experienced users with top of line hardware and current software. Here is an example:

• 75Mhz Pentium Packard Bell, 850MB,8MB Memory/1MB Video/Quad Speed CD,14.4 Fax Modem/Speakers/spkr phone,14” SVGA Color Monitor.

• Linux Red Hat Beginner Pack,Windows ’95, Encarta ’95, Fax, MS Money, Action Pres, and much more!

• Customization: Dual boot between Linux and Windows ’95, Home users desktop ready and customized, Your Internet provider access preped for Windows ’95 and Linux Slip*, and other features to meet your needs.

Estimated Cost: $1988 Packard Bell and Windows Software + $139 Linux Beginer and Integration = $2127 total. Will spend 15-30 minutes over the phone to acquire the machine requirements as you wish. For short time, Dr. Linux Book included at no cost with the machine integration offering! 3 year parts and labor warranty for $80. Beat that! (first year on-site!). Smaller/other machines available (486 and 60Mhz Pentiums, hundreds less, similar software features)! Finally, we provide Advanced User Updates:

• quarterly: $22.50/Quarter

• bi-quarterly: $28.50/Semi Quarterly

• yearly: $36.50/Yearly

We offer customer discounts on other Oxbow Distribution products based upon update rate. Quarterly: 10% Semi-quarterly: 7% Each update includes 10 e-mail queries and 30 minutes of technical support.

Ordering:

Check or Money Order, Visa and MC in near future. Machine purchase and non-linux materials purchased for you in your name. Shipping extra. Prices effective through 10/15/95. Machine purchases require money orders at the current time.

Entry last modified:

23 November 1995

5 Miscellany

This section lists distributors selling commercial software for Linux (such as Motif) and miscellany, such as documentation.

5.1 Unifix Custom Linux CD-ROM Mastering

Distributor:

Unifix Software GmbH, Braunschweig, Germany Postfach 4918 D-38039 Braunschweig, Germany Vox: +49 (0)531 515161 Fax: +49 (0)531 515162

Provider’s Description:

We use our own (Linux based) premastering system to make custom CDs which include the customer’s programs in addition to our Linux distribution. This enables our customers to sell their Unix based applications together with operating system tested and ready to run. Our CD recorder makes it possible to create prototypes and very small series at interesting prices.

Ordering:

Ask us for more information or prices.

5.2 Linux Journal

Publisher:

Linux Journal P.O. Box 85867, Seattle, WA 98145-1867 Net: ljeditor@ssc.com Vox: (206) 782-7733 (206-PUBS-REF) Fax: 206-782-7191

Provider’s Description:

Linux Journal is a monthly publication covering the Linux Community. Most material in LJ is new (not reprinted from Usenet). Each issue includes columns and articles on Linux programming, Free Software Foundation issues, systems administration, Questions and Answers, interviews and more. LJ is a professional-quality magazine for the Linux community.

Ordering:

Subscriptions are $22/year (US), $27/year (Canada/Mexico), $32/year (elsewhere). Payment can be made using VISA, MasterCard or American Express (make sure you include the credit card number, expiration date and signature). We can also accept checks in US $ on US banks. If you have a question concerning appropriate method of payment, phone or FAX our subscription numbers or send e-mail to subs@ssc.com For security reasons we discourage sending credit card numbers via e-mail. You can get a PGP key at <info@ssc.com>.

Miscellaneous:

If you are interested in advertising in LJ, call, write or e-mail at linux@ssc.com. Article queries, new product announcements and other editorial material should be sent to our address above or e-mailed

to ljeditor@sunsite.unc.edu. Generic questions can be sent to linux@ssc.com.

Entry last modified:

24 May 1996.

5.3 Sequoia International Motif Development Package

Distributor:

Sequoia International, Inc. 600 West Hillsboro Blvd, Suite 300, Deerfield Beach, FL 33441, Net: info@seq.com Vox: (305) 480-6118 Fax: (305) 480-6198

Provider’s Description:

Sequoia International, Inc. has a complete Motif 1.2.3 Runtime and Development package called SWiM 1.2.3 available for $149.95. In addition to providing shared library versions of libXm and libMrm, the following is included in each package: The Window Manager (mwm), Shared Libraries (libXm, libMrm), (libXm, libMrm, libUil), UIL Compiler, Header and Include Files, Complete On-Line Manual Pages, Source Code to OSF/Motif Demo Programs, Complete OSF/Motif Users Guide.

Requirements:

Linux 0.99pl13 or higher, libc 4.4.4, XFree 2.0, 12M Free Disk, 8-12M RAM Suggested.

Ordering:

USA: Sequoia International, Inc. (305-480-6118), info@seq.com. Japan: Fortune Co., Ltd (03-5481-8974). England: Lasermoon Ltd. (+44-0-329-826444). Australia: Space Age Import-Export Proprietary (61-7-266-3418).

Entry last modified:

24 March 1994.

5.4 SSC Books and Linux Documentation Project Manuals

Distributor:

SSC, Inc. P.O. Box 55549 Seattle, WA 98155 Net: sales@ssc.com Vox: 206-782-7191 (206-PUBS-REF) Fax: 206-782-7191

Provider’s Description:

SSC has printed the Linux Installation and Getting Started manual, Version 2.2.2, to make it available for those who do not have the capability to print it themself and to support our customers who buy Linux distributions from us. The book is perfect-bound with a 2-color cover. SSC has also published the Network Administrator’s Guide from the LDP and will soon publish the Kernel Hacker’s Guide. In addition, SSC has published The Linux Sampler, a book which includes new material plus articles from back issues of Linux Journal. Other products include Linux T-shirts, mouse pads and bumper stickers.

Ordering:

Linux Installation and Getting Started, Version 2.1 is available for $12.95 plus shipping ($3 in the U.S.). The Linux Network Administrator’s Guide and The Linux Sampler are $18.95 each plus shipping. Finger bookshelf@ssc.com for information on these and other Linux books that we will make available. We can accept credit card orders (Visa, MasterCard or AmEx). Orders can be phoned in, faxed, mailed, or emailed.

Miscellaneous:

SSC also publishes a series of reference cards on Unix and Unix-related programs such as Emacs, VI, Korn Shell, C language, etc. SSC also sells the InfoMagic, Yggdrasil and Trans Ameritech distributions, Free Software Foundation books and a complete Linux package consisting of Linux Instllation and Getting Started, a Linux distribution on CD, 4 SSC Pocket References and a 1-year subscription to Linux Journal for $64.95. Call or e-mail sales@ssc.com for a free catalog.

Entry last modified:

28 May 1996.

5.5 Extent Verlag, LDP Distribution

Distributor:

Extent Verlag Berlin, Germany Postfach 12 66 48, D-10594 Berlin, Germany Vox: +49 30 3244021 Fax: +49 30 3249685

Provider’s Description:

Extent has published the Linux Installation and Getting Started manual, Version 2.0 by Matt Welsh in order to make it available to Linux users within Germany and Europe. Inspite of being non-commercial, the manual was printed with 2540 dpi and bound in soft cover. Its handy paperback format is 148 x 210 mm. Other LDP manuals are being prepared as soon as possible. Extent also tries to make SLACKWARE available to everybody within Germany and Europe just for the ordinary diskette price. This means, you pay only the diskettes and get the newest SLACKWARE distribution for free, just in the sense of Free Software. The 3.5" diskettes are fully error-checked, of course.

Ordering:

Linux Installation and Getting Started, 192 pages. ISBN 3-926671-12-2. Costs within Germany DM 15.80 (tax and shipping included), within Europe DM 16.50 (shipping included). Full Slackware (latest version) on 50 diskettes. Within Germany, DM 100.00 plus DM 10.00 for shipping. Within Europe, DM 100.00 plus DM 20.00 for shipping. Ordering by sending cheque to address above or money order (don’t forget your address!) to Extent Verlag Berlin, Germany Postbank Berlin, BLZ 10010010, bank account 1769-104 or every book store.

Entry last modified:

25 March 1994.

5.6 UPython SLS and Slackware Diskette Labels

Distributor:

UPython Computer and Network Services

Provider’s Description:

Labels for Softlanding and slackware disks

Internet Access:

The labels that we print for you are available in several formats at

<sunsite.unc.edu:/pub/linux/distributions/SLSlabels>.

Ordering:

For order information, send mail to DGray@uh.edu. Cost is $6.50 per set and domestic shipping in included. Overseas orders must send payment in US funds and add $1 for shipping. Texas residents add sales tax (6%).

Entry last modified:

11 Nov 1993.

6 Submissions To This Document

I encourage anyone and everyone who maintains a Linux distribution or mail order service, to submit information on their service to this HOWTO. It’s easy and fun, and it’s free advertising. This document is posted to many places and is archived (see the next section).

6.1 Types of Submissions

We are interested in submissions for:

• Complete distributions of Linux software, available either via anonymous FTP, UUCP, or mail order. By ‘complete distribution’ we mean any set of software which can be used to build a complete Linux system from scratch.

• Layered products or individual software packages available only via mail order. If your software package is available via anonymous FTP, chances are people can find it. Software products only available via mail-order include commercial things such as Motif, and any other commercial software ported to Linux.

• Any other Linux-specific goods available via mail order, such as Linux-specific books and documentation, T-shirts, and assorted paraphrenalia. Note: our interest in the latter categories is marginal – this is a Linux Distributions HOWTO, not a catalogue of cutesy Linux junk. Whatever you’re selling will have to be (a) very useful, (b) very original, or (c) very funny, to get listed here.

If the number of submissions for services and layered products is large, I’ll create a separate HOWTO for these items.

6.2 How to submit

To submit an entry to this HOWTO, please send mail to esr@snark.thyrsus.com with the following information. This format is not machine-parsable; any of the fields may be any length that you wish, but I’d like to keep each entry down to, say, 50 lines.

Name:

Name of service or distribution Distributor:

Name of company, person, etc. who distributes/maintains the service or distribution. Should include mail, email, phone contact information, and (if possible) a WWW reference.

Provider’s Description:

Description of the distribution or service that you provide. If this is a software distribution, please include information such as what software is included, versions, general overview of installation, requirements, and so on.

Internet Access:

Where your service or distribution is available over the Internet; typically a WWW or FTP address.

Ordering:

How to order your distribution or service, if applicable. Include prices, shipping information, methods of payment, etc.

Support:

Information on support terms and support contracts.

Miscellaneous:

Anything else that you find relevant.

Last Freeze Date:

Last freeze date of the current version(s). Also, your estimated update frequency.

6.3 Submission Guidelines

Please keep your entry as short as possible. If you need to include extensive information, please make a reference to where one can FTP or mail to get more information on your distribution; these entries are only meant to be pointers to where one can find information on your service or distribution.

If you provide more than one service or distribution, please use separate entries for each.

Some things (such as books, t-shirts, etc.) won’t fit exactly into this format; just be sure to include all relevant information. In other words, this ”entry form” is simply the bare minimum that you must include in your submission to the Distribution-HOWTO; feel free to change, add, or leave out ”fields” as you see fit.

I may edit your entries for conciseness and brevity, if I find any irrelevant information, or if the entry is overly verbose. Otherwise the content should remain the same.

When making submissions to the Distribution-HOWTO, you grant implicit permission for me to use the entries in other materials, such as books from the LDP, and other online documents. For example, information from the Distribution-HOWTO may be included in a published Linux book. If you do not want me to include your entry in materials other than the Distribution-HOWTO, please say so.

7 Administrivia

7.1 Terms of Use

This document is copyright 1996 by Eric S. Raymond. You may use, disseminate, and reproduce it freely, provided you:

• Do not omit or alter this copyright notice.

• Do not omit or alter or omit the version number and date.

• Do not omit or alter the document’s pointer to the current WWW version.

• Clearly mark any consdensed, altered or versions as such.

These restrictions are intended to protect potential readers from stale or mangled versions. If you think you have a good case for an exception, ask me.

7.2 Acknowledgements

This document was originated by Bill Riemers. Matt Welsh maintained the second version. Erik Troan (ewt) then maintained the document to release 3.0.

In January 1995, Eric Raymond, while unaware of the existence of this document, began to develop a similar FAQ in the format of his now-discontinued ”PC-clone UNIX Software Buyer’s Guide”, which had covered mainly System V UNIXes and BSD/OS.

In March 1995, Eric approached Erik about cooperating on a merged version. In early April 1995, Erik went to work for Red Hat Software, and (wishing to avoid a conflict of interest) handed the document to Eric. Eric merged in a lot of new information and added several new fields to the distribution entries.

Accordingly, this document has been a sort of serial collaboration. The editorial ‘we’ generally tags observations by all the maintainers; ‘I’ is Eric (the current one) speaking.

We are delighted to acknowledge the contributions of all the Linux users and Internet hackers who have contributed information and feedback.